# Evolving Dark Sector and the Dark Dimension Scenario

### v1

@@ Abstract
String theory naturally leads to the expectation that dark energy is not stable, and may be evolving as captured by the Swampland de Sitter conjectures. Moreover, motivated by the distance conjecture a unification of dark sector has been proposed, where the smallness of dark energy leads to one extra dimension of micron size with dark matter being the Kaluza–Klein graviton excitations in this extra dimension. In this paper we consider the natural possibility that the radius of the dark dimension varies as the dark energy decreases, leading to the variation of the dark matter mass. This correlates the decrease of the dark energy with the variation of the dark matter mass as they depend on the variations of a scalar field $\phi$ controlling the radius of the extra dimension. A simple realization of this idea for small range of $\phi$ is adequately captured by choosing a potential which is locally of the form $V=V_{0}\ {\rm exp}(-c\phi)$ and dark matter mass $m_{\rm DM}=m_{0}\ {\rm exp}(-c^{\prime}\phi)$ where the sign of $\phi$ is chosen such that $c^{\prime}\geq 0$ while we have two choices for the sign of $c$ depending on whether the dark dimension expands or shrinks when the dark energy dominates. We find that this scenario is in excellent agreement with recent experimental data from DESI DR2 combined with SN measurements (from DES, Union3 or PantheonPlus) and reproduces the same significance as CPL parametrization with the added benefit of providing a natural explanation for the apparent phantom behavior ($w<-1$) reported by DESI and DES based on a physical model. Regardless of the SN dataset, there is a preference for non-zero values of $c^{\prime}$ and $c$ that are in the expected $O(1)$ range in Planck units as suggested by the Swampland criteria. In particular, there is a remarkable consistency with $c^{\prime}\simeq 0.05\pm 0.01$ for all dataset combinations including SN, and close to the experimental upper bound of $c^{\prime}\lesssim 0.2$ demanded by the lack of detection of fifth force in the dark sector.
@@ I Introduction
In non-gravitational theories, a common paradigm that guides model building is the concept of distance scale separation, which is the idea that microscopic physics has limited influence on macroscopic phenomena. This leads to the idea that for an EFT in the infrared (IR) one can ignore the ultraviolet (UV) description as all the terms in the action are allowed subject to symmetries of the theory where the strengths of the higher derivative terms are set by the UV cutoff. This principle underlies many cosmological models that aim to describe the universe on the largest scales. However, in gravitational theories, it has been understood that UV and IR physics are deeply interconnected. The Swampland program (see Agmon et al. (2022); van Beest et al. (2022); Palti (2019) for reviews), rooted in lessons learned from string theory, seeks to make the IR implications of a consistent gravitational UV completion more precise. In particular, the form of the scalar potential and the spectrum of light states—both crucial to the cosmological history of the universe—are tightly constrained by the requirement of UV completeness in quantum gravity.
@@
In Agrawal et al. (2018), it was shown that the dS conjecture is compatible with cosmological observations. Moreover, motivated by these conjectures, Swampland compatible models of evolving dark energy have been considered in many works. For example, in Agrawal et al. (2021) it was pointed out that the natural unification of the dark energy and dark sector as implied by the Swampland criteria would lead to the fading of the dark matter (decreasing of $\rho_{M}a^{3})$ as the universe evolves and the dark energy depletes. The main aim of this paper is to reevaluate this proposal in light of the dark dimension scenario and the recent DES and DESI DR2 results. In particular, we view the evolving length of the dark dimension as the mechanism that lowers the dark energy and depending on whether it expands or shrinks it decreases or increases the dark matter mass. Indeed, we find that this scenario is in excellent agreement with data, with preference for shrinking dark dimension in recent epoch which reproduces the same significance as the CPL parametrization with the benefit of providing a natural explanation for the apparent phantom behavior ($w<-1$) reported by DESI and DES.
@@ II.1 Scalar potential
[…]
@@
[…]
@@ II.2 Tower of light states
[…]
@@ III Dark Dimension Scenario
with $m\sim 1/L$ and for some order 1 parameter $A$, which leads to a micron scale extra dimension. Moreover, it was suggested that the associated light tower, the KK graviton in the 5-th dimension, serves as a natural Gonzalo et al. (2023) and viable candidate Obied et al. (2024); Law-Smith et al. (2024) for dark matter (see Vafa (2024) for a review). This is a dynamical model for dark matter Dienes and Thomas (2013), where there is intra-tower KK graviton decays which on its own does not change the total mass of the dark matter due to an approximate conservation of KK number which follows from the smoothness of the dark dimension. Note, however, the total mass of the dark matter sector can also change if the length of the dark dimension changes, as the tower states have masses roughly of the order of $n/L$ for integer spaced quanta $n$. Indeed, we will consider this possibility in this paper (for another approach to dark matter in dark dimensions see Anchordoqui et al. (2023a, b, 2022, 2024)).
@@ IV Application to Cosmology
The directions of $\nabla m$ and $\nabla V$ do not need to align. In particular $c^{\prime}$ can naturally be smaller than $|\nabla m|/m$ if we view $\phi$ as parameterizing $\nabla V$ direction and if $\nabla m$ is not aligned with it. Moreover, if their inner product is negative, the projection of these vectors onto the single-field trajectory that drives the cosmological evolution yields a negative value for $c$. This situation can naturally arise in the interior of field space, as illustrated in Fig. 2, where the scalar potential possesses a local maximum. If the scalar field lies to the left of this maximum, then $\nabla m$ and $\nabla V$ point in opposite directions. It is important to note, however, that in such a case the field will eventually roll towards the asymptotic region $\phi\rightarrow-\infty$, where a different tower of states becomes the lightest.
@@
Components of the energy density driving the expansion of our universe are often modeled as perfect fluids. These are characterized by an equation of state parameter $w$ which is the ratio of the fluid pressure to its energy density. In general, the equation of state is time-dependent although it is constant for many common fluids like cold dark matter, non-relativistic baryons, massless radiation, and the cosmological constant. The relation between energy density and expansion scale depends on the fluid type with the relation $\rho\propto a^{-3(1+w)}$. The cosmological constant, which has $w=-1$ and an energy density that does not dilute with the expansion of the universe, marks the limiting case of this behavior. No known fluids have $w<-1$ which would mean that the energy density increases with the expansion of the universe. This behavior violates the null energy condition and would require a more exotic fluid known as phantom.
@@
In light of the above theoretical discussion, it is interesting that recent measurements made by the DESI collaboration (and analyzed in conjunction with CMB and supernova data and the recent results of DES) show evidence of an equation of state for dark energy that is less than $-1$ Lodha et al. (2025) and thus violates the null energy condition.
@@
The key to understanding the phantom behavior is that the dark energy equation of state cannot be defined unambiguously in the presence of a DM/DE coupling. This is because, when such a coupling exists, the interaction term contributes to the energy momentum tensor. Moreover, the individual energy densities are no longer simple functions of the scale factor, and the naive definition $w=-\frac{1}{3}\frac{d\log(\rho_{\rm DE})}{d\log(a)}$ is influenced by the interaction. In particular, if one assumes that DM redshifts as $w=0$ (i.e., like $1/a^{3}$), then one can subtract its contribution from the total energy density and fit the evolution of DE accordingly. In the presence of interactions between DM and DE, this effective equation of state is not constrained to satisfy $w>-1$, and in fact, for certain couplings, we can obtain $w<-1$. Since the new DESI data analysis assumes $\rho_{\rm DM}\sim 1/a^{3}$, we advocate that the evidence for $w_{\rm DE}<-1$ found by the DESI collaboration is actually evidence for a coupling between dark energy and dark matter.
It is easy to understand intuitively why a coupling between dark matter and dark energy can lead to an effective dark energy equation of state $w_{\mathrm{DE,eff}}<-1$, and we will see that this indeed occurs in our model. Consider the evolution of the combined DM + DE fluid in $\Lambda$CDM, where the total energy density redshifts in a prescribed manner. Now, suppose there is a coupling between DM and DE that causes this combined fluid to dilute more slowly than in $\Lambda$CDM. Heuristically, this can occur if DM transfers part of its energy density to DE, although, as noted earlier, it is not possible to define these energy densities separately and precisely in the presence of interactions. Nevertheless, if one models this interacting system under the assumption that DM still redshifts as $1/a^{3}$, then one is forced to infer the existence of a DE component with $w<-1$, since this is the only way to achieve slower dilution than in $\Lambda$CDM while keeping the matter dilution rate and density at early times fixed.
@@
which satisfies $-1\leq w_{\phi}\leq 1$ with the limiting cases corresponding to a scalar field with vanishing kinetic energy and potential respectively. In term of the quantity $w_{\phi}$, we write the effective DE equation of state as:
@@
For a non-rolling scalar field, we have $\phi=\phi_{i}$ and $w_{\rm{DE,eff}}\rightarrow w_{\phi}$ satisfying the usual expectation that uncoupled quintessence DE has an equation of state between $-1$ and $1$. On the other hand, if the scalar field rolls to larger values as the universe expands, i.e. $\phi_{i}<\phi$, then we have that $x<0$ and $w_{\rm{DE,eff}}<w_{\phi}$ which allows $w_{\rm{DE,eff}}<-1$ if $w_{\phi}$ is sufficiently close to $-1$. This is indeed the behaviour we get in the fading DM model. The equations of motion for the scale factor and the scalar field are
@@
The fact that $w_{\rm{DE,eff}}$ can have a magnification factor compared to $w_{\phi}$ and make it appear phantom-like was already noted in Stefancic (2004); Huey and Wandelt (2006); Das et al. (2006); van de Bruck and Thomas (2019); Agrawal et al. (2021); Khoury et al. (2025). However, except for Agrawal et al. (2021), which is the model we are considering here, all the other works assumed the mass of the dark matter increases as the scalar potential decreases, so that the effective potential acquires a minimum. Another key difference is in the definition of the effective dark energy density. Except for Agrawal et al. (2021), to attain a phantom behavior despite the fact that mass goes up, they define an effective dark energy density by subtracting the redshifted contribution of the current dark matter density instead of the initial one [^4] Since the sampling of the data is scattered across different epochs the effective observational $w_{\rm{DE,eff}}$ is presumably a subtraction of fixed dark matter density at some intermediate point.. The combination of these two differences leads for those models to potentially also have an effective phantom behavior.
@@
Figure 3: Contours showing the 1- and 2-$\sigma$ constraints on $c$ and $c^{\prime}$, along with their marginalized posteriors for two fading dark sector models with opposite signs of $c$, across various datasets. $\Lambda$CDM corresponds to $c=c^{\prime}=0$.
@@
Given that our model is similar to the fading dark matter model of Agrawal et al. (2021) and that earlier results there show its consistency with the best-fit $w_{0}$$w_{a}$ parameterization, we expect that it can reproduce the phantom-like behavior preferred by the DESI+CMB(+SN) datasets within a well-motivated physical framework. To explore this, we re-analyze the fading dark matter model using the newly available data. Specifically, we run Metropolis–Hastings Markov Chain Monte Carlo (MCMC) simulations using the Boltzmann solver CLASS Blas et al. (2011) and the sampler Cobaya Torrado and Lewis (2019, 2021). The MCMCs are performed with the following datasets:
  • Baryon Acoustic Oscillations (BAO) measurements from DESI DR2 Abdul Karim et al. (2025), which provide measurements of the transverse comoving distance $D_{M}(z)=\int_{0}^{z}\rm{d}z^{\prime}\,H(z^{\prime})^{-1}$ and the line-of-sight Hubble distance $D_{H}(z)=H(z)^{-1}$ for various tracers.
@@
$\phi$$\tau$$m_{\rm DM}$$z$
(a) $c<0,c^{\prime}>0$
$\phi$$\tau$$m_{\rm DM}$$z$
(b) $c>0,c^{\prime}>0$
Figure 5: Evolution of the scalar field and normalized mass scale of the tower that makes up the dark matter for two fading dark sector models with opposite signs of $c$, shown across various datasets. In both cases, the dark matter mass exhibits a net decrease. However, for $c<0$, a brief increase occurs during dark energy domination due to the reversal in direction of the scalar field that controls the dark matter mass.
@@
DESI+CMB $-12.0$   $3.0$ $-8.8$   $2.5$ $-1.03$ $0.05$ $-6.8$   $2.1$ $0.42$ $0.05$
$-0.72^{+0.45}_{-0.43}$ $0.05\pm 0.01$ $0.43^{+0.28}_{-0.29}$ $0.05\pm 0.01$
DESI+CMB+Union3 $-16.8$   $3.7$ $-14.6$   $3.4$ $-1.12$ $0.05$ $-10.9$   $2.9$ $0.81$ $0.06$
$-1.01^{+0.21}_{-0.22}$ $0.05\pm 0.01$ $0.71^{+0.22}_{-0.24}$ $0.05\pm 0.01$
DESI+CMB+Pantheon+ $-10.2$   $2.7$ $-11.5$   $2.9$ $-0.85$ $0.05$ $-9.4$   $2.6$ $0.65$ $0.05$
$-0.76^{+0.23}_{-0.22}$ $0.05\pm 0.01$ $0.54^{+0.21}_{-0.24}$ $0.05\pm 0.01$
DESI+CMB+DESY5 $-20.3$   $4.1$ $-19.4$   $4.0$ $-1.06$ $0.05$ $-15.9$   $3.6$ $0.83$ $0.06$
$-1.00^{+0.13}_{-0.14}$ $0.05\pm 0.01$ $0.79^{+0.14}_{-0.12}$ $0.05\pm 0.01$
Table 1: Statistical significance comparing the fading dark sector models for both $c>0$ and $c<0$ alongside the CPL parametrization, across various combinations of supernova datasets, to $\Lambda$CDM. Additionally, the best fit, the mean, and the ±$\pm 1\sigma$ values for $c,c^{\prime}$ are provided for each fading dark sector model.
@@
Given that the scalar field $\phi$ is very light, with Hubble mass, one might worry about the implications of a fifth force mediated by $\phi$. Let us make two important remarks. First, since visible matter is localized on the Standard Model brane in the extra dimension, it does not couple directly to $\phi$. As a result, no fifth force is mediated within the visible sector. Second, regarding the dark sector, bounds on fifth forces between dark matter particles that rely on standard cosmological evolution do not directly apply to our model, which exhibits a different cosmological history. There are also bounds on fifth forces from galactic-scale observations Kesden and Kamionkowski (2006); Mohapi et al. (2016) that must be taken into account.
@@
$\frac{F_{\phi}}{F_{\rm grav}}=2c^{\prime 2}\,.$ (18)
Observational bounds constrain the ratio between the dark force and the gravitational force in the dark sector to be less than an $\mathcal{O}(1)$ number. At first glance, this may appear to require fine-tuning, especially given that gravity is the weakest of the fundamental forces. However, according to (18), the Swampland Distance Conjecture suggests that the expected strength of a scalar-mediated force relative to gravity for a tower of particles is naturally of $\mathcal{O}(1)$, thereby resolving this apparent fine-tuning problem. For instance, bounds derived from orbital dynamics, such as those in Kesden and Kamionkowski (2006); Mohapi et al. (2016), suggest $c^{\prime}\lesssim 0.2$. Yet the precise upper limit remains model-dependent, since these bounds rely on orbital periods spanning billions of years and assume standard $\Lambda$CDM cosmology. Modifications such as the evolution and coupling of dark matter and dark energy, as realized in our model, can alter these constraints. Therefore, it would be interesting and worthwhile to revisit such analyses within our dynamical DM/DE framework. It is also reassuring that the best-fit value from our model, $c^{\prime}\simeq 0.05\pm 0.01$, satisfies the current upper bound inferred from Sagittarius stream data within $\Lambda$CDM. The fact that the upper bound is only slightly above the observed value of $c^{\prime}$ in our model raises the exciting prospect of detecting a dark fifth force in the near future.

### v2

@@ Abstract
String theory naturally leads to the expectation that dark energy is not stable, and may be evolving as captured by the Swampland de Sitter conjectures. Moreover, motivated by the distance conjecture, a unification of dark sector has been proposed, where the smallness of dark energy leads to one extra dimension of micron size with dark matter being the Kaluza–Klein graviton excitations in this extra dimension. We consider the natural possibility that the radius of the dark dimension varies as the dark energy decreases, leading to the variation of the dark matter mass. This correlates the decrease of the dark energy with the variation of the dark matter mass as they depend on the variations of a scalar field $\phi$ controlling the radius of the extra dimension. A simple realization of this idea for small range of $\phi$ is adequately captured by choosing a potential which is locally of the form $V=V_{0}\ {\rm exp}(-c\phi)$ and dark matter mass $m_{\rm DM}=m_{0}\ {\rm exp}(-c^{\prime}\phi)$ where the sign of $\phi$ is chosen such that $c^{\prime}\geq 0$ while we have two choices for the sign of $c$ depending on whether the dark dimension expands or shrinks when the dark energy dominates. We find excellent agreement with recent experimental data from DESI DR2 combined with SN measurements (from DES, Union3 or Pantheon+) and reproduces the same significance as CPL parametrization with the added benefit of providing a natural explanation for the apparent phantom behavior (w<1w<-1) reported by DESI and DES based on a physical model. DESI and SN datasets independently favor non-zero values of $c^{\prime}$ and $c$, respectively, both lying within the expected $\mathcal{O}(1)$ range suggested by the Swampland criteria. Moreover, our best fit value $c^{\prime}\simeq 0.05\pm 0.01$ is remarkably consistent with the experimental upper bound of $c^{\prime}\lesssim 0.2$ demanded by the lack of detection of fifth force in the dark sector.
@@ I Introduction
In non-gravitational theories, a common paradigm that guides model building is the concept of distance scale separation, which is the idea that microscopic physics has limited influence on macroscopic phenomena. This leads to the idea that for an EFT in the infrared (IR) one can ignore the ultraviolet (UV) description as all the allowed terms in the action are subject to symmetries of the theory where the strengths of the higher derivative terms are set by the UV cutoff. This principle underlies many cosmological models that aim to describe the universe on the largest scales. However, in gravitational theories, it has been understood that UV and IR physics are deeply interconnected. The Swampland program (see Agmon et al. (2022); van Beest et al. (2022); Palti (2019) for reviews), rooted in lessons learned from string theory, seeks to make the IR implications of a consistent gravitational UV completion more precise. In particular, the form of the scalar potential and the spectrum of light states—both crucial to the cosmological history of the universe—are tightly constrained by the requirement of UV completeness in quantum gravity.
@@
In Agrawal et al. (2018), it was shown that the dS conjecture is compatible with cosmological observations. Moreover, motivated by these conjectures, Swampland compatible models of evolving dark energy have been considered in many works. For example, in Agrawal et al. (2021) it was pointed out that the natural unification of the dark energy and dark sector as implied by the Swampland criteria would lead to the fading of the dark matter (decreasing of $\rho_{M}a^{3})$ as the universe evolves and the dark energy depletes. The main aim of this paper is to reevaluate this proposal in light of the dark dimension scenario and the recent DES and DESI DR2 results. In particular, we view the evolving length of the dark dimension as the mechanism that lowers the dark energy and depending on whether it expands or shrinks it decreases or increases the dark matter mass. Indeed, we find that this scenario is in excellent agreement with data, with preference for shrinking dark dimension in recent epoch which reproduces the same significance as the CPL parametrization with the benefit of providing a natural explanation for the apparent phantom behavior (w<1w<-1) reported by DESI and DES.
@@ II.1 Scalar potential
$V(\phi)\sim\exp(-\lambda\phi)\quad\Rightarrow\quad\lambda\geq\frac{2}{\sqrt{d-2}}=\sqrt{2}\quad{\it for}\ d=4\,.$ (3)
@@
$\Rightarrow\lambda_{\rm interior}\geq\sqrt{\frac{2}{3}}\approx 0.82\quad{\it for}\ d=4\,.$ (4)
@@ II.2 Tower of light states
$\sqrt{\frac{d-1}{d-2}}\geq\lambda^{\prime}\geq\frac{1}{\sqrt{d-2}}\Rightarrow\sqrt{\frac{3}{2}}\geq\lambda^{\prime}\geq\frac{1}{\sqrt{2}}\quad{\it for}\ d=4\,.$ (6)
@@ III Dark Dimension Scenario
with $m\sim 1/L$ and for some order 1 parameter $A$, which leads to a micron scale extra dimension. Moreover, it was suggested that the associated light tower, the KK graviton in the 5-th dimension, serves as a natural Gonzalo et al. (2023) and viable candidate Obied et al. (2024); Law-Smith et al. (2024) for dark matter (see Vafa (2024) for a review). This is a dynamical model for dark matter Dienes and Thomas (2012a); *Dienes:2011sa. In the scenario of Gonzalo et al. (2023) there is intra-tower KK graviton decays which on its own does not appreciably change the total mass of the dark matter due to an approximate conservation of KK number which follows from the smoothness of the dark dimension. Note, however, the total mass of the dark matter sector can change if the length of the dark dimension changes, as the tower states have masses roughly of the order of $n/L$ for integer spaced quanta $n$. Indeed, we will consider this possibility in this paper (for another approach to dark matter in dark dimensions see Anchordoqui et al. (2023a, b, 2022, 2024)).
@@ IV Application to Cosmology
The directions of $\nabla m$ and $\nabla V$ do not need to align. In the dark dimension scenario, $m$ is only a function of the size of the fifth dimension. Assuming that there is only one light scalar field $\phi$ which is parallel to $\nabla V$, depending on the change in the size of the fifth dimension in that direction, $c^{\prime}$ can be much smaller than 1 in Planck units [^3] As seen in Fig. S10, data does not prefer a cosmology driven by a more complicated 2D trajectory in the scalar field space.. For example, if changing $\phi$ leaves the diameter of the fifth dimension intact, $c^{\prime}$ would vanish. However, a more natural possibility is that $c^{\prime}$ is $\mathcal{O}(1)$ in Planck units indicating that dark dimension diameter varies as $\phi$ flows. We choose the direction of $\phi$ such that $m$ decreases as $\phi$ increases. Note that if the inner product $\nabla m\cdot\nabla V$ is negative, the single-field trajectory parametrized by $\phi$ yields a negative value for $c$. This situation can naturally arise in the interior of field space, as illustrated in Fig. 2, where the scalar potential possesses a local maximum. If the scalar field lies to the left of this maximum, then $\nabla m$ and $\nabla V$ point in opposite directions.
@@
Let us comment on how non-zero values of $c$ and $c^{\prime}$ affect different epochs of cosmology. During the radiation-dominated era, the scalar field remains effectively frozen due to Hubble friction being set by the radiation energy density. As a result, the dynamics are largely unaffected in this epoch. However, during the matter-dominated era, the situation changes: as the scalar field rolls, the dark matter density decreases due to the exponential dependence of the mass $m(\phi)$ on $\phi$. Consequently, this epoch and any observations probing it (e.g. DESI) will be most sensitive to the parameter $c^{\prime}$. On the other hand, the dark energy–dominated era is more sensitive to modifications of the scalar potential itself and therefore provides a probe of $c$. In this epoch, the scalar field rolls in the direction of $\rm{sgn}(c)$ so that the dark energy density decreases due to the exponential dependence of the scalar potential $V(\phi)$ on $\phi$. Therefore, the supernovae datasets which probe low redshifts should be expected to be most sensitive to $c$. Thus, one can intuitively associate the dominant effects of $c^{\prime}$ and $c$ with distinct cosmological epochs: the matter-dominated era and the dark energy–dominated era, respectively. For non-zero values of $c$ and $c^{\prime}$, we expect the dark matter to fade in the matter-dominated era and dark energy to fade in the dark energy–dominated era, which is why we call this model the Fading Dark Sector model. Note that if $c$ and $c^{\prime}$ have opposite signs, the dark energy increases in the dark matter dominated era and the dark matter increases in the dark energy–dominated era which is a secondary effect in the expansion of the universe.
@@
Components of the energy density driving the expansion of our universe are often modeled as perfect fluids. These are characterized by an equation of state parameter $w$ which is the ratio of the fluid pressure to its energy density. In general, the equation of state is time-dependent although it is constant for many common fluids like cold dark matter, non-relativistic baryons, massless radiation, and the cosmological constant. The relation between energy density and expansion scale depends on the fluid type with the relation $\rho\propto a^{-3(1+w)}$. The cosmological constant, which has w=1w=-1 and an energy density that does not dilute with the expansion of the universe, marks the limiting case of this behavior. No known fluids have w<1w<-1 which would mean that the energy density increases with the expansion of the universe. This behavior violates the null energy condition and would require a more exotic fluid known as phantom.
@@
In light of the above theoretical discussion, it is interesting that recent measurements made by the DESI collaboration (and analyzed in conjunction with CMB and supernova data and the recent results of DES) show evidence of an equation of state for dark energy that is less than 1-1 Lodha et al. (2025); Gialamas et al. (2025) and thus violates the null energy condition.
@@
The key to understanding the phantom behavior is that the dark energy equation of state cannot be defined unambiguously in the presence of a DM/DE coupling. This is because, when such a coupling exists, the interaction term contributes to the energy momentum tensor. Moreover, the individual energy densities are no longer simple functions of the scale factor, and the naive definition $w=-\frac{1}{3}\frac{d\log(\rho_{\rm DE})}{d\log(a)}$ is influenced by the interaction. In particular, if one assumes that DM redshifts as $w=0$ (i.e., like $1/a^{3}$), then one can subtract its contribution from the total energy density and fit the evolution of DE accordingly. In the presence of interactions between DM and DE, this effective equation of state is not constrained to satisfy w>1w>-1, and in fact, for certain couplings, we can obtain w<1w<-1. Since the new DESI data analysis assumes $\rho_{\rm DM}\sim 1/a^{3}$, we advocate that the evidence for $w_{\rm DE}<-1$ found by the DESI collaboration is actually evidence for a coupling between dark energy and dark matter.
[…]
@@
which satisfies $1\leq w_{\phi}\leq 1$ with the limiting cases corresponding to a scalar field with vanishing kinetic energy and potential respectively. In term of the quantity $w_{\phi}$, we write the effective DE equation of state as:
@@
For a non-rolling scalar field, we have $\phi=\phi_{i}$ and $w_{\rm{DE,eff}}\rightarrow w_{\phi}$ satisfying the usual expectation that uncoupled quintessence DE has an equation of state between 1-1 and $1$. On the other hand, if the scalar field rolls to larger values as the universe expands, i.e. $\phi_{i}<\phi$, then we have that $x<0$ and $w_{\rm{DE,eff}}<w_{\phi}$ which allows $w_{\rm{DE,eff}}<-1$ if $w_{\phi}$ is sufficiently close to 1-1. This is indeed the behaviour we get in the fading DM model. The equations of motion for the scale factor and the scalar field are
@@
The fact that $w_{\rm{DE,eff}}$ can have a magnification factor compared to $w_{\phi}$ and make it appear phantom-like was already noted in Stefancic (2004); Huey and Wandelt (2006); Das et al. (2006); van de Bruck and Thomas (2019); Agrawal et al. (2021); Chakraborty et al. (2025); Khoury et al. (2025). However, except for Agrawal et al. (2021), which is the model we are considering here, all the other works assumed the mass of the dark matter increases as the scalar potential decreases, so that the effective potential acquires a minimum. Another key difference is in the definition of the effective dark energy density. Except for Agrawal et al. (2021), to attain a phantom behavior despite the fact that mass goes up, they define an effective dark energy density by subtracting the redshifted contribution of the current dark matter density instead of the initial one [^5] Since the sampling of the data is scattered across different epochs the effective observational $w_{\rm{DE,eff}}$ is presumably a subtraction of fixed dark matter density at some intermediate point.. The combination of these two differences leads for those models to potentially also have an effective phantom behavior.
@@
[…]
@@
Given that our model is similar to the fading dark matter model of Agrawal et al. (2021) and that earlier results there show its consistency with the best-fit $w_{0}$$w_{a}$ parameterization, we expect that it can reproduce the phantom-like behavior preferred by the DESI+CMB(+SN) datasets within a well-motivated physical framework. To explore this, we re-analyze the fading dark matter model using the newly available data. In order to run Markov Chain Monte Carlo (MCMC) simulations, we adapt the publicly available Boltzmann solver CLASS Blas et al. (2011) to include scalar fields coupled to a single species of Dark Matter (DM). We modify the evolution of the background cosmology and also include the new (coupled) perturbations in a self-consistent manner so that our modeling fully takes into account the effect of perturbations on linear scales. The couplings, potentials, and dependence of the DM mass on the scalar fields (in the background and perturbation modules) are implemented in a way that is easy to modify and we use this freedom to test various models (such as those discussed in the Supplemental Materials Supplemental Figures and Tables).
We use the Metropolis–Hastings algorithm built into the sampler Cobaya Torrado and Lewis (2019, 2021) to sample from the parameter posteriors in each case. Our MCMCs are run including the following datasets:
  • Baryon Acoustic Oscillations (BAO) measurements from DESI DR2 Abdul Karim et al. (2025), which provide measurements of the transverse comoving distance $D_{M}(z)=\int_{0}^{z}\mathrm{d}z^{\prime}\,H(z^{\prime})^{-1}$ and the line-of-sight Hubble distance $D_{H}(z)=H(z)^{-1}$ for various tracers.
@@
$\phi$$\tau$
(a) Scalar field trajectory.
$m_{\rm DM}$$z$
(b) Dark matter mass.
$V_{\rm DE}$$z$
(c) Dark energy density.
Figure 5: Evolution of the scalar field and normalized mass scale of the tower that makes up the dark matter for two fading dark sector models with opposite signs of $c$, shown across various datasets. Here, the solid and dashed lines indicate the $c<0$ and $c>0$ FDS model, respectively. In both cases, the dark matter mass exhibits a net decrease. However, for $c<0$, there is a brief increase during dark energy domination due to the reversal in direction of the scalar field that controls the dark matter mass. We also show the potential of the dark energy for both of the FDS models.
@@
DESI+CMB 12.0-12.0 $3.0$ 8.8-8.8 $2.5$ 1.03-1.03 $0.05$ 6.8-6.8 $2.1$ $0.42$ $0.05$
$72^{+0.45}_{-0.43} 0.0$$05\pm 0.01$ $43^{+0.28}_{-0.29} 0.0$$05\pm 0.01$
DESI+CMB+Union3 16.8-16.8 $3.7$ 14.6-14.6 $3.4$ 1.12-1.12 $0.05$ 10.9-10.9 $2.9$ $0.81$ $0.06$
$1.01^{+0.21}_{-0.22}$ $0.05\pm 0.01$ $71^{+0.22}_{-0.24} 0.0$$05\pm 0.01$
DESI+CMB+Pantheon+ 10.2-10.2 $2.7$ 11.5-11.5 $2.9$ 0.85-0.85 $0.05$ 9.4-9.4 $2.6$ $0.65$ $0.05$
$76^{+0.23}_{-0.22} 0.0$$05\pm 0.01$ $54^{+0.21}_{-0.24} 0.0$$05\pm 0.01$
DESI+CMB+DESY5 20.3-20.3 $4.1$ 19.4-19.4 $4.0$ 1.06-1.06 $0.05$ 15.9-15.9 $3.6$ $0.83$ $0.06$
$1.00^{+0.13}_{-0.14}$ $0.05\pm 0.01$ $79^{+0.14}_{-0.12} 0.0$$05\pm 0.01$
[…]
@@
As we explained in Section IV, experiments probing the dark matter–dominated epoch, such as DESI which has most of its data from that era, primarily constrain the parameter $c^{\prime}$, while low-redshift measurements, such as supernovae datasets that are most sensitive to the dark energy–dominated era, serve as probes of $c$. These two types of observations are nearly independent. For instance, excluding the supernovae datasets leads to the same best-fit value of $c^{\prime}$ with $c$ compatible with 0 (see Table 1), whereas removing DESI and including only supernovae data selects the same value of $c$ but yields a different best-fit value of $c^{\prime}$ compatible with 0 (see Table 2). Moreover, the significances of the individual fits approximately add up to the significance of the combined fit in Table 1. This indicates that interpreting $c$ and $c^{\prime}$ as independent variables constrained by independent datasets is a good approximation.
One might wonder how much of our analysis is sensitive to the shape of the scalar potential. For example, let us consider replacing the scalar potential with a hilltop potential in the Fading Dark Sector model. Since DESI prefers a non-zero value of $c^{\prime}$, the scalar field rolls significantly during the matter-dominated era regardless of the shape of the scalar potential. Therefore, unless we unnaturally fine-tune the initial condition for $\phi$ such that the scalar field lands at the top of the hilltop at the end of the matter dominated era, we are only sensitive to the slope of the scalar potential in the dark energy dominated era. In particular, as seen in Fig. S6, if we choose the initial value of the scalar field to be at the top of the hilltop, it rolls off the hilltop before dark energy dominates, leading to the same picture with an exponential potential with $c>0$. This agreement confirms that only the local slope of the potential captured by $c$ is sufficient to parameterize $V$ as we have assumed. Adding spatial curvature to the Fading Dark Sector does not improve the fit either as data prefers no spatial curvature (Figs. S8, S9). Alternatively, if one were to replace the fading dark matter with spatial curvature (as proposed in Akrami et al. (2025)) the result would be statistically disfavored relative to the Fading Dark Sector (Table 3). Our results differ from those in Akrami et al. (2025) for runs with $\Omega_{k}\neq 0$ due to our use of the full CMB dataset and inclusion of SN samples beyond DESY5, which notably worsen the fits.
Given that the scalar field $\phi$ is very light, with Hubble mass, one might worry about the implications of a fifth force mediated by $\phi$. Let us make two important remarks. First, since visible matter is localized on the Standard Model brane in the extra dimension, it does not couple directly to $\phi$. As a result, no fifth force is mediated within the visible sector. Second, regarding the dark sector, bounds on fifth forces between dark matter particles that rely on standard cosmological evolution do not directly apply to our model, which exhibits a different cosmological history. There are also bounds on fifth forces from galactic-scale observations Kesden and Kamionkowski (2006) that must be taken into account.
@@
$\frac{F_{\phi}}{F_{\rm grav}}=2c^{\prime 2}\simeq 0.005\pm 0.002\,.$ (18)
The fifth force in the dark sector is strongly constrained to be small. The smallness of the fifth dark force is naturally explained in our model given that $c^{\prime}$ is $\mathcal{O}(1)$ according to (18), thereby resolving this apparent fine-tuning problem. In particular, the bound derived in Kesden and Kamionkowski (2006) from tidal disruption on dwarf galaxies orbiting the Milky Way suggests $c^{\prime}\lesssim 0.2$. It is reassuring that the best-fit value from our model, $c^{\prime}\simeq 0.05\pm 0.01$, corresponding to an attractive fifth force in the dark sector $0.5\%$ of the strength of the gravitational force, satisfies the current upper bound. The fact that the upper bound is only slightly above the observed value of $c^{\prime}$ in our model raises the exciting prospect of detecting a dark fifth force in the near future. It would be interesting and worthwhile to revisit the bounds on the fifth force within our dynamical DM/DE framework.
Our best-fit value for the fifth-force ratio, $F_{\phi}/F_{\rm grav}\simeq 0.005$, is consistent with the analysis of Bottaro et al. (2025), which independently found a hint for a 5-th force in the dark sector $F_{5}/F_{\rm grav}\simeq 0.004\pm 0.002$ in the DESI DR1 data [^7] We thank Mikhail Ivanov and Tracy Slatyer for bringing this paper to our attention after the first version of this paper appeared..